OpenStudy (anonymous):

Help please! (: Solve each equation by undoing each operation: 2(x+4)^2-7=0

jimthompson5910 (jim_thompson5910):

I'll get you started \[\large 2(x+4)^2-7=0\] \[\large 2(x+4)^2=7\] \[\large (x+4)^2=\frac{7}{2}\]

jimthompson5910 (jim_thompson5910):

tell me what you get when you finish that up

OpenStudy (anonymous):

do u divide 7 and 2?

jimthompson5910 (jim_thompson5910):

no leave that as a fraction if you want an exact answer

jimthompson5910 (jim_thompson5910):

your next step is to undo the square

OpenStudy (anonymous):

Im not sure how to do that:/ sorry! would you get 4x^2=4/2 or would you square root both sides..

jimthompson5910 (jim_thompson5910):

you would apply the square root to both sides

jimthompson5910 (jim_thompson5910):

since the square root undoes the square

OpenStudy (anonymous):

mmkay, so: \[4+ or - \sqrt{7/2}\]

jimthompson5910 (jim_thompson5910):

close

OpenStudy (anonymous):

im confused..

jimthompson5910 (jim_thompson5910):

well you had x+4 so to undo the +4, you subtract 4 from both sides

jimthompson5910 (jim_thompson5910):

which means the 4 should be negative

OpenStudy (anonymous):

oh, okay. Thanks!

jimthompson5910 (jim_thompson5910):

and keep in mind that \[\large \sqrt{\frac{7}{2}} = \frac{\sqrt{7}}{\sqrt{2}}\] \[\large \sqrt{\frac{7}{2}} = \frac{\sqrt{7}\sqrt{2}}{\sqrt{2}\sqrt{2}}\] \[\large \sqrt{\frac{7}{2}} = \frac{\sqrt{7*2}}{\sqrt{2*2}}\] \[\large \sqrt{\frac{7}{2}} = \frac{\sqrt{14}}{\sqrt{4}}\] \[\large \sqrt{\frac{7}{2}} = \frac{\sqrt{14}}{2}\]

OpenStudy (anonymous):

Oh my goodness, haha(: thanks!!

jimthompson5910 (jim_thompson5910):

so \[\large x = -4 \pm \sqrt{\frac{7}{2}}\] turns into \[\large x = -4\pm\frac{\sqrt{14}}{2}\]

jimthompson5910 (jim_thompson5910):

and finally... \[\large x=-4\pm\frac{\sqrt{14}}{2}\] \[\large x=-\frac{8}{2}\pm\frac{\sqrt{14}}{2}\] \[\large x=\frac{-8\pm\sqrt{14}}{2}\]

jimthompson5910 (jim_thompson5910):

technically once you've isolated x, you're done...but...some books will want it in certain forms and this is usually a form you'll see the answer in

OpenStudy (anonymous):

haha(: wow! thanx! GBU!!

jimthompson5910 (jim_thompson5910):

yw